OpenStudy (anonymous):

find the sum enter your answer in lowest terms 6/17 + 2/8

OpenStudy (anonymous):

\[\large\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\] and don't let anyone tell you different

OpenStudy (anonymous):

Find common numerator, 68, and solve for answers (hint, 2/8= 1/4)

OpenStudy (anonymous):

in this case you get \[\large \frac{6\times 8+17\times 2}{17\times 8}\] that is all

OpenStudy (anonymous):

there is no such thing as a "common numerator"

OpenStudy (abb0t):

LOL

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

doesn't anyone own a calculator any more? one with a fraction key?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

really? even on a computer? of course you do

OpenStudy (anonymous):

no

OpenStudy (anonymous):

 what is 17 x 8 ?